OpenStudy (anonymous):

Evaluate: 5(6 − 3) − 1 • 6

OpenStudy (anonymous):

fo the parentheses and product. 5.3-6=15-6=9

OpenStudy (anonymous):

so the answer is 9

OpenStudy (anonymous):

yes.